Building A was built 20 years before building B and 30 years after building C. In 10 years, building C will be 20 years younger than the combined ages of buildings A and B. What is the present age of each building?
I started to set up the problem and solve it, but my solution doesn't make any sense.
BUILDING A : Now = x-20 -> In 10 yrs = (x-20)+10
BUILDING B : Now = x -> In 10 yrs = x+10
BUILDING C : Now = x+30 -> In 10 yrs = (x+30)+10
BUILDING B
[(x+30)+10]-20=(x-20)+10+x+10
x+20=2x -> x-x+20=2x-x
x=20 <- Doesn’t Make Any Sense for the Age of B Now.

Answer by stanbon(75887)

Building A was built 20 years before building B and 30 years after building C. In 10 years, building C will be 20 years younger than the combined ages of buildings A and B. What is the present age of each building?
-----------
Let age of B be "x".
Then age of A is "x+20"
And age of C is "x + 30"
---
Equation:
x+30+10 = (x+10)+(x+20+10)-20
---
x + 40 = 2x+20
x = 20 age of B
x + 20 age of A
x + 30 age of C
=====================
Cheers,
Stan H.
==============